# IGEV++: Iterative Multi-range Geometry Encoding Volumes for Stereo Matching

### v1

@@ IV-G Real-Time Version of IGEV++
To meet the requirement of time-constrained applications, we further demonstrate that the proposed IGEV++ can be easily configured to achieve real-time inference on KITTI-resolution ($1248\times 384$) images. Compared to IGEV++, the real-time version (named RT-IGEV++), is based on four main modifications. First, we remove the context network and use a single backbone (feature network) to generate context features. Second, we construct a single-range GEV with a pre-defined maximum disparity of 192px instead of MGEV. Third, since our GEV already integrates contextual cues and contains non-global geometry information, we use a single-level ConvGRU to replace RAFT-Stereo’s three-level ConvGRUs. Finally, for efficiency, we reduce the hidden state channels of the ConvGRU from 128 to 96. As shown in Table VII, our RT-IGEV++ achieves the best performance among all published methods on the Scene Flow test set. We present results for iteration numbers of 2, 4, 6, and 8, enabling users to trade off between time efficiency and performance according to their needs. We also evaluate RT-IGEV++ on the KITTI 2012 and 2015 test sets, the results are shown in Table III. Our RT-IGEV++ ranks first among all published real-time methods. On KITTI 2012, RT-IGEV++ outperforms HITNet [17] and Fast-ACVNet+ [27] by 8.51% and 11.03% on 3-noc metric, respectively. On KITTI 2015, RT-IGEV++ surpasses HITNet [17] and Fast-ACVNet+ [27] by 13.64% and 10.95% on D1-all metric, respectively.
@@ IV-H Zero-shot Generalization Performance
Since large-scale real-world datasets for training models are difficult to obtain, the zero-shot generalization ability of stereo models is crucial. We evaluate the zero-shot generalization performance of IGEV++ from synthetic Scene Flow datasets to unseen real-world scenes. In this evaluation, we train our IGEV++ on the Scene Flow dataset [14], and then directly test it on the Middlebury training sets. As shown in Table VIII, our IGEV++ achieves state-of-the-art performance in the zero-shot setting. Figure 2 shows the visual comparisons with PCWNet [11], DLNR [16], GMStereo [5], our IGEV++ effectively distinguishes subtle details in complex backgrounds.
@@ V Conclusion
To efficiently handle large disparities and textureless/reflective regions, we further propose multi-range geometry encoding volumes (MGEV) that encode coarse-grained geometry information for textureless regions and large disparities and fine-grained geometry information for details and small disparities. To effectively and efficiently construct MGEV and fuse geometry features across multiple ranges and granularities within MGEV, we introduce an adaptive patch matching module and a selective geometry feature fusion module, respectively.

### v2

@@ IV-G Real-Time Version of IGEV++
To meet the requirement of time-constrained applications, we further demonstrate that the proposed IGEV++ can be easily configured to achieve real-time inference on KITTI-resolution ($1248\times 384$) images. Compared to IGEV++, the real-time version (named RT-IGEV), is based on four main modifications. First, we remove the context network and use a single backbone (feature network) to generate context features. Second, we construct a single-range GEV with a pre-defined maximum disparity of 192px instead of MGEV. Third, since our GEV already integrates contextual cues and contains non-global geometry information, we use a single-level ConvGRU to replace RAFT-Stereo’s three-level ConvGRUs. Finally, for efficiency, we reduce the hidden state channels of the ConvGRU from 128 to 96. As shown in Table VIII, our RT-IGEV achieves the best performance among all published methods on the Scene Flow test set. We present results for iteration numbers of 2, 4, 6, and 8, enabling users to trade-off between time efficiency and performance according to their needs. We also evaluate RT-IGEV on the KITTI 2012 and 2015 test sets, the results are shown in Table IV. Our RT-IGEV ranks first among all published real-time methods. On KITTI 2012, RT-IGEV outperforms HITNet [17] and Fast-ACVNet+ [27] by 8.51% and 11.03% on 3-noc metric, respectively. On KITTI 2015, RT-IGEV surpasses HITNet [17] and Fast-ACVNet+ [27] by 13.64% and 10.95% on D1-all metric, respectively.
@@ IV-H Zero-shot Generalization Performance
Since large-scale real-world datasets for training models are difficult to obtain, the zero-shot generalization ability of stereo models is crucial. We evaluate the zero-shot generalization performance of IGEV++ from synthetic Scene Flow datasets to unseen real-world scenes. In this evaluation, we train our IGEV++ on the Scene Flow dataset [14], and then directly test it on the Middlebury V3 training sets. As shown in Table IX, our IGEV++ achieves state-of-the-art performance in the zero-shot setting. Figure 2 shows the visual comparisons with PCWNet [11], DLNR [16], GMStereo [5], our IGEV++ effectively distinguishes subtle details in complex backgrounds.
@@ V Conclusion
To efficiently handle large disparities and textureless/reflective regions, we further propose multi-range geometry encoding volumes (MGEV), which encode coarse-grained geometry information for ill-posed regions and large disparities, while preserving fine-grained geometry information for details and small disparities. To effectively and efficiently construct MGEV and fuse geometry features across multiple ranges and granularities within MGEV, we introduce an adaptive patch matching module and a selective geometry feature fusion module, respectively.
@@
We also propose RT-IGEV, a real-time version of IGEV++, which achieves real-time inference while delivering the best performance among all published real-time methods.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (623B2036, 62472184) and the Fundamental Research Funds for the Central Universities.